# Demystifying Invariant Effectiveness for Securing Smart Contracts

### v1

@@ 6.2. RQ2: Study of False Positives and True Positives
We also manually analyze the false positives generated by our invariants to assess their impact on regular, benign users. For this, we sample up to 10 transactions from the false positives for each invariant and manually evaluate their bypassability. We operate under the assumption that regular users have the option to split transactions with large parameters into transactions with smaller parameters (via DApp’s front-end), lower the gas for their transactions (via a wallet software), or simply wait for some time to transact again after their transaction previously blocked. Based on these criteria, we categorize the false positives into three groups: D1: the transaction is completely blocked and cannot be bypassed through simple means; D2: the transaction can be bypassed by breaking it down into smaller transactions; and D3: users can bypass the transaction by reducing the gas or waiting for some time. For both true positives and false positives, two authors independently labeled the bypassability results with the third one to resolve the divergence of views.
@@
For certain invariants such as TOU and DFU, there is a dichotomy where they block exploit transactions effectively but normal users can still bypass them. They make certain exploits impossible by preventing hackers from reaching specific contract states required for profitability. Regular users, who usually do not attempt to manipulate contract states for exploitative gains, can often bypass these invariants by simply splitting their larger transactions into smaller ones. Though this results in more transactions and higher gas costs, it does not impede their primary objectives.
@@ 6.3. RQ3: Effectiveness of Combination of Invariants
Experiment. In RQ3, we look into the effectiveness of various invariant combinations in safeguarding smart contracts. Building on insights from RQ1 and RQ2, we identify a set of five effective invariants (EOA, SO, GC, TOU, and DFU), replacing GS with GC considering that it can be easily bypassed. Our aim is to explore whether these invariants, each effective in its own domain, can be combined to provide a more robust defense against contract exploits while maintaining a low false positive rate.
@@ Results.
Average FP rate(%) 3.99 D1 8 Average FP rate(%) 0.32 D1 10
@@
The combination EOA $\land$ (OB $\lor$ DFU) maintains an impressively low false positive rate of 0.32%, making it superior based on the metric (2). Its efficacy is due in part to OB’s inherent low false positive rate and to the complementary nature of OB and DFU in the context of token transfer functions. Both combined invariants are better at stopping hacks than individual invariants. This is because they can work together in different parts of the smart contract code, making it more likely they will catch harmful actions. Sometimes a function in the contract may only be protected by one invariant, if other invariants are not applicable to this function. In those cases, both hackers and normal users may still find a way to bypass the security measures.
Answer to RQ3: The invariant guards studied and generated by Trace2Inv are complimentary and their combinations are promising to be more effective on contract protection and hack prevention with lower false positive rate.
@@ 6.4. RQ4: Gas Overhead of Invariant Guards
Experiment. In our next research question, RQ4, we examine the gas overhead incurred by the deployment of individual and combined invariants. We select four benchmark contracts that represent different kinds of protocols and programming languages. The target smart contract is instrumented with the generated invariants studied in RQ3 and then compiled by either a Solidity or Vyper compiler. This process is carried out for both the original and the instrumented versions of the contract. The compiler returns an estimate of the gas consumption for each function in the contract, allowing us to calculate the gas overhead for each inserted invariant by comparing the two versions. Subsequently, we replay all transactions within the test set on these instrumented contracts. For each transaction, we add the corresponding gas overhead when the transaction reaches where the invariants are inserted.
@@
Bridge HarmonyBridge Solidity 0.5.17 0.04 0.52 0.02 0.01 0.07 0.56
Lending Harvest1 Solidity 0.5.18 0 1.46 0.01 0.01 0.02 1.46
Yield-Earning CreamFi2_1 Solidity 0.5.17 0 2.55 0 0 0.01 2.55
Others BeanstalkFarms Vyper 0.2.8 0 0.53 / 0 0 0.53
@@
Answer to RQ4: The gas overheads of these invariant guards are as low as 0% - 2.55%.
@@ 6.5. RQ5: Comparative Analysis with Other State-of-the-art (SOTA) Tools
Experiment 1: Compare with InvCon+, a SOTA invariant mining tool. InvCon+ (Liu et al., 2024), a direct follow-up work of InvCon (Liu and Li, 2022), leverages transaction pre/post conditions to generate invariants aimed at mitigating real-world smart contract vulnerabilities. InvCon only infers likely invariants and only produces raw results from Daikon (Ernst et al., 2007). In contrast, InvCon+ generates accurate invariants that are verified against the contract’s transaction history. Similar to Trace2Inv, InvCon+ takes a target contract and its transaction history as inputs and automatically generates invariants. We contacted authors of InvCon+ and obtained the InvCon+ tool. We then applied it on our benchmarks in Table 4. Following the same methodology of RQ1 and RQ3, we used $70\%$ of the transaction history for training, testing the generated invariants on the remaining $30\%$.
Experiment 2: Compare with TxSpector, a SOTA transaction attack detection tool. Though Trace2Inv is not primarily designed for transaction anomaly detection, we explored its ability to flag attack transactions. TxSpector (Zhang et al., 2020), a SOTA framework for transaction attack detection, identifies attacks using eight detectors: re-entrancy, unchecked call, failed send, timestamp dependence, unsecured balance, misuse of origin, suicidal, and gas-related re-entrancy. TxSpector takes a transaction as input and automatically flags it if it is identified as an attack. We obtained TxSpector from its public repository [^1] https://github.com/OSUSecLab/TxSpector, and applied it to the testing set of our benchmarks. Same as what is used in TxSpector paper (Zhang et al., 2020), we assigned a timeout of $60$ seconds for all benign transactions. For hack transactions which are usually more complex, we assigned a timeout of $2$ hours.
@@
# Hacks(27) 20 Blocked 6 Blocked 7 Detected
@@
In contrast, Trace2Inv, utilizing the invariant template EOA $\land$ (OB $\lor$ DFU), effectively secured 31 out of 42 victim contracts across 20 hacks with a remarkably low false positive rate of just 0.32%. Moreover, Trace2Inv demonstrated enhanced practicality by generating an average of only $12$ invariants per contract, significantly outperforming InvCon+ in terms of real-world viability. Unlike TxSpector, which only detects hacks, Trace2Inv not only blocks a greater number of hacks but also achieves this with a significantly lower false positive rate, indicating its effectiveness in anomaly detection as well.
Identifying New Exploits: In the development of Trace2Inv, we surprisingly found two previously unreported exploit transactions, earlier than any reported exploit transactions against RariCapital1222https://etherscan.io/tx/0x4764dc6ff19a64fc1b0e57e735661f64d97bc1c44e026317be8765358d0a7392 and Yearn, [^3] https://etherscan.io/tx/0x59faab5a1911618064f1ffa1e4649d85c99cfd9f0d64dcebbc1af7d7630da98b respectively. These two transactions were initially reported by Trace2Inv as false positives, because they were not flagged as attacks when we collected benchmarks. However, after manual investigation, we found that these two transactions caused a huge financial loss and the addresses of the originators of the two transactions are flagged on EtherScan as “Rari Capital Exploiter” and “Yearn (yDai) Exploiter”, respectively. Thus, we believe these are indeed exploit transactions. This discovery further underscores Trace2Inv’s potential in unveiling new exploits.
@@ 7. Discussion
This study contributes valuable insights into the application and effectiveness of invariant-based security measures across different types of decentralized finance (DeFi) protocols. In this section, we discuss some key takeaways and their implications for future research and practical applications.
@@
Choosing Complementary Invariants. Our findings underscore the importance of selecting a diverse set of invariants to safeguard smart contracts. Each type of invariant serves as a unique line of defense against abnormal transactions. For instance, invariants in time lock category act as temporal barriers, making it difficult for attackers to execute key functions like withdraw multiple times within a single transaction. On the other hand, invariants in data flow and money flow categories limit the token amounts that can be withdrawn in one function call. By employing a combination of these invariants, developers force attackers to only withdraw a controlled amount of tokens per transaction, which may disrupt the underlying logic that the attack relies on, thereby blocking the attack.
@@
Dynamic Parameter Updates for Invariants. Another key insight from our research is the need for dynamic parameter updates for certain types of invariants. For invariants that are tied to variables that change cumulatively over time—like oracle prices or storage values—parameters in the invariants can quickly become obsolete. If such an invariant is violated, it could lead to a cascade of failed transactions, causing a high False Positive rate. Therefore, it is crucial for developers to continuously monitor these variables and adjust the invariant parameters. Conversely, for invariants related to independent actions like token transfers, the parameters can remain relatively stable, as user behavior in these domains tends to be stable over time.
@@
Mitigating Flash Loan Attacks. Our benchmarks indicate a significant prevalence of flash loan-based exploits, with 17 out of 27 examined exploits leveraging flash loan. Flash loans enable users to borrow large amounts of tokens for the duration of a single transaction, providing attackers with substantial resources to execute complex exploits. Our approach of enforcing invariants can effectively mitigate the risk posed by flash loans. Many effective invariants such as EOA and OB block flash loan attacks by enforcing attackers to split their transaction into multiple ones. Without flash loan, the attacker would need to use their own assets to execute the attack with the uncertainty of other bots’ backrunning between the hackers’ transactions. This raise not only the technical but also financial barriers to successful attacks.
@@
The Impact of Invariants on Contract Composability. Incorporating invariant guards into smart contracts might limit their adaptability and integration with DeFi protocols. However, our study in Section 4 reveals that many invariants stem from existing DeFi protocol requirements, underscoring the preference of developers for security benefits over flexibility. Moreover, our findings in RQ2, as discussed in Section 6, show that certain invariants, like OB, has almost no false positives, while others, such as TOU and DFU, can possibly block malicious transactions while allowing normal users to circumvent them. These insights imply that with careful selection and understanding of target user behaviors, developers can devise invariant guards that minimally affect contract composability.
@@ 8. Related Work
Smart Contract Invariants. There exist several works studying smart contract invariants (Liu et al., 2022a, b; Deng et al., 2024; Zhou et al., 2020). Zhou et al. (Zhou et al., 2020) introduces $6$ invariants to defend against different hacks. Cider (Liu et al., 2022a) leverages deep reinforcement learning on smart contract source code to learn invariants that prevent arithmetic overflows. SPCon (Liu et al., 2022b) utilizes function callers in past transactions of a contract to recover a likely access control model. More recently, Over (Deng et al., 2024) infers safety constraints on oracles from the contract’s source code and the history of oracle updates. Compared to the above works, Trace2Inv studies a broader range of invariants and evaluates their effectiveness and bypassability against real-world attacks.
@@
Smart Contract Security Analysis. There are a large body of works on detection of smart contract security vulnerabilities (Feist et al., 2019; Brent et al., 2020; Tikhomirov et al., 2018; Sofware Reliability Lab, 2019; Feng et al., 2019; Luu et al., 2016; man, 2019; myt, 2019; Kalra et al., 2018; Jiang et al., 2018; Wang et al., 2020; Wüstholz and Christakis, 2020; Trail of Bits, 2019). Oyente (Luu et al., 2016) is one of the earliest symbolic execution-based security tools to detect reentrancy, mishandled exception, transaction order dependence, and timestamp dependence. It has been extended to detect greedy, prodigal and suicidal contracts (Nikolić et al., 2018). Other well-known symbolic-execution tools also include Manticore (man, 2019) and Mythril (myt, 2019) which are able to find other types of vulnerabilities, e.g., dangerous delegatecall, integer overflow, etc. Slither (sli, 2021) is another popular static security analysis tool for smart contracts. It performs data flow and control flow dependency analysis to support up to 87 bug detectors including dangerous strict equality and the dangerous usage of tx.origin. Other static analyzers include SmartCheck (Tikhomirov et al., 2018) mainly targeting bad coding practices, Securify (Sofware Reliability Lab, 2019) and Ethainter (Brent et al., 2020) for finding information-flow vulnerabilities. Moreover, many dynamic analysis tools (Jiang et al., 2018; Atzei et al., 2017; Nguyen et al., 2020; Trail of Bits, 2019; Wüstholz and Christakis, 2020; Wang et al., 2020; Liu et al., 2020) were proposed to detect smart contract vulnerabilities through fuzzing and model-based testing.
@@
There is a large body of work on the detection and exploitation of DeFi vulnerabilities (Wu et al., 2021; Kong et al., 2023; Qin et al., 2021; Baum et al., 2022; Xue et al., 2022; Deng et al., 2023; Gudgeon et al., 2020). Qin et al. (Qin et al., 2021) and Zhou et al. (Zhou et al., 2021) formulated oracle price manipulation and arbitrage as an optimization problem to identify attack opportunities maximizing attacker’s profit. Wu et al. (Wu et al., 2021) identified several oracle price manipulation patterns from on-chain transaction data to detect real-world attacks while Kong et al. (Kong et al., 2023) detects price manipulation vulnerabilities in DeFi applications through inter-contract taint analysis. Also, Gudgeon et al. (Gudgeon et al., 2020) showcased how to explore flashloan to conduct governance attack. Baum et al. (Baum et al., 2022) surveyed the state-of-the-art mitigation techniques for front-running in DeFi, such as fair ordering, batching of blind inputs, private and secrete state. Interestingly, several works (Xue et al., 2022; Deng et al., 2023; Zhang et al., 2023) explored front-running as a defense mechanism against smart contract exploits. Different from the over-generalized security patterns used by the existing tools, our invariant guards capture the subtle semantic constraints of specific smart contracts.
@@
Runtime Verification and Validation. Runtime verification is a light-weight verification approach that monitors the operation of a running system, detects and possibly reacts to unsatisfied observed behaviors against certain properties. Runtime verification can be used for validation purpose where the runtime checks are on properties from users’ expectation rather than from formal program specifications (Magazzeni et al., 2017). Sereum (Rodler et al., 2018) is a general runtime validation framework to protect deployed contracts against reentrancy attacks. Sereum extends Ethereum by introducing the detection module for monitoring attacks rooted in different types of reentrancy. Solythesis (Li et al., 2020) provides a source-to-source compiler that facilitates the runtime validation of smart contracts. Specifically, Solythesis takes as inputs a smart contract code and a user specified invariant. It generates an enhanced smart contracts that reject all unexpected contract transactions. Additionally, the overhead of the runtime validation approach is shown to be quite small. Trace2Inv distinguishes itself from the above-mentioned approaches in that the existing approaches only offer limited protection for several fixed types of attacks. In contrast, Trace2Inv is not bound to specific attack types and has been proven effective in mitigating many sophisticated attack vectors.

### v2

@@ 6.2. RQ2: Study of False Positives and True Positives
We also manually analyze the false positives generated by our invariants to assess their impact on regular users. For this, we sample up to 10 transactions from the false positives for each invariant and evaluate their bypassability. We operate under the assumption that regular users have the option to split transactions with large parameters into transactions with smaller parameters, lower the gas for their transactions, or simply wait for some time to transact again after their transaction is blocked. Based on these criteria, we categorize the false positives into three groups: D1: transaction is completely blocked and cannot be bypassed through simple means; D2: transaction can be bypassed by breaking it down into smaller transactions; and D3: users can bypass the transaction by reducing the gas or waiting for some time. For both true and false positives, two authors independently labeled the bypassability results with the third author to resolve divergence of views.
@@
For certain invariants such as TOU and DFU, there is a dichotomy where they block exploit transactions effectively, but normal users can still bypass them. They make certain exploits impossible by preventing hackers from reaching specific contract states required for profitability. Regular users, who usually do not attempt to manipulate contract states for exploitative gains, can often bypass these invariants by simply splitting their larger transactions into smaller ones. Though this results in more transactions and higher gas costs, it does not impede their primary objectives.
@@ 6.3. RQ3: Effectiveness of Combination of Invariants
Experiment. In RQ3, we look into the effectiveness of various invariant combinations in safeguarding smart contracts. Building on insights from RQ1 and RQ2, we identify a set of five effective invariants (EOA, SO, GC, TOU, and DFU), replacing GS with GC considering that it can be easily bypassed. Our aim is to explore whether these invariants, each effective in its own domain, can be combined to provide a more robust defense against hacks while maintaining a low FP rate.
@@ Results.
Average FP rate(%) 3.99 D1 8 Average FP rate(%) 0.28 D1 10
@@
The combination EOA $\land$ (OB $\lor$ DFU) maintains an impressively low false positive rate of 0.28%, making it superior based on the metric (2). Its efficacy is due in part to OB’s inherent low false positive rate and to the complementary nature of OB and DFU in the context of token transfer functions. Both combined invariants are better at stopping hacks than individual invariants. This is because they can work together in different parts of the smart contract code, making it more likely they will catch harmful actions. Sometimes a function in the contract may only be protected by one invariant, if other invariants are not applicable to this function. In those cases, both hackers and normal users may still find a way to bypass the security measures.
Answer to RQ3: The invariant guards studied and generated by Trace2Inv are complimentary and their combinations are promising to be more effective on contract protection and hack prevention with lower false positive rates.
@@ 6.4. RQ4: Gas Overhead of Invariant Guards
Experiment. In RQ4, we examine the gas overhead incurred by the deployment of individual and combined invariants. We select four benchmark contracts that represent different kinds of protocols and programming languages. The target smart contract is instrumented with the generated invariants studied in RQ3 and then compiled by either a Solidity or Vyper compiler. This process is carried out for both the original and the instrumented versions of the contract. The compiler returns an estimate of the gas consumption for each function in the contract, allowing us to calculate the gas overhead for each inserted invariant by comparing the two versions. Subsequently, we replay all transactions within the test set on these instrumented contracts. For each transaction, we add the corresponding gas overhead when the transaction reaches where the invariants are inserted.
@@
Bridge HarmonyBridge Solidity 0.5.17 0.04 0.59 $\rightarrow$ 0.12+ 0.02 0.01 0.07 0.63 $\rightarrow$ 0.16
Lending Harvest1 Solidity 0.5.18 0.00 0.96 $\rightarrow$ 0.02 0.01 0.01 0.02 0.97 $\rightarrow$ 0.04
Yield-Earning CreamFi2_1 Solidity 0.5.17 0.00 0.14 $\rightarrow$ 0.01 0.01 0.00 0.01 0.15 $\rightarrow$ 0.01
Others BeanstalkFarms Vyper 0.2.8 0.00 4.09 $\rightarrow$ 0.68 / 0.00 0.00 4.09 $\rightarrow$ 0.68
@@
Answer to RQ4: The gas overheads of these invariant guards are as low as 0% - 0.68%.
@@ 6.5. RQ5: Comparative Analysis with Other State-of-the-art (SOTA) Tools
Experiment 1: Compare with InvCon+, a SOTA invariant mining tool. InvCon+ (Liu et al., 2024), a direct follow-up work of InvCon (Liu and Li, 2022), leverages transaction pre/post conditions to generate invariants aimed at mitigating real-world smart contract vulnerabilities. InvCon only infers likely invariants and only produces raw results from Daikon (Ernst et al., 2007). In contrast, InvCon+ generates accurate invariants that are verified against the contract’s transaction history. Similar to Trace2Inv, InvCon+ takes a target contract and its transaction history as inputs and automatically generates invariants. We obtained InvCon+ (inv, 2024) from its authors, and applied it on our benchmarks in Table 4. Following the same methodology of RQ1 and RQ3, we used $70\%$ of the transaction history for training, testing the generated invariants on the remaining $30\%$.
Experiment 2: Compare with TxSpector, a SOTA transaction attack detection tool. Though Trace2Inv is not primarily designed for transaction anomaly detection, we explored its capability to identify attack transactions. TxSpector (Zhang et al., 2020) identifies attacks using eight detectors: re-entrancy, unchecked call, failed send, timestamp dependency, unsecured balance, misuse of origin, suicidal, and gas-related re-entrancy. We obtained TxSpector from its public repository (txs, 2020). TxSpector operates with a modified Geth archive node to produce transaction traces that are fed into the detectors. Using components of Trace2Inv along with an unmodified Geth archive node, we generated equivalent transaction traces for the testing sets of our benchmarks, and applied TxSpector detectors to analyze them. We also patch TxSpector to support new EVM opcodes introduced after its last update. Consistent with the setup in the TxSpector paper (Zhang et al., 2020), we set a timeout of $60$ seconds for all benign transactions. We extended this to $2$ hours for hack transactions, which are typically more complex, when applying the detectors.
@@
# Hacks(27) 20 Blocked 3 Blocked 7 Detected
@@
In contrast, Trace2Inv, utilizing the invariant template EOA $\land$ (OB $\lor$ DFU), effectively secured $31$ of $42$ victim contracts across $20$ hacks with a remarkably low FP rate of just $0.28$%. Moreover, Trace2Inv demonstrated enhanced practicality by generating an average of only $12$ invariants per contract, significantly outperforming InvCon+ in terms of real-world viability. Unlike TxSpector, which only detects hacks, Trace2Inv not only blocks a greater number of hacks but also achieves this with a significantly lower FP rate, indicating its effectiveness in anomaly detection as well.
Identifying New Exploits: In the development of Trace2Inv, we surprisingly found two previously unreported exploit transactions, earlier than any reported exploit transactions against RariCapital1 (Rar, 2023a) and Yearn (Yea, 2024). The two transactions were initially reported by Trace2Inv as false positives, because they were not flagged as hacks when we collected benchmarks. However, after manual investigation, we found the two transactions caused a huge financial loss and the addresses of the originators of the two transactions are flagged on EtherScan as “Rari Capital Exploiter” and “Yearn (yDai) Exploiter”, respectively. Thus, we believe they are indeed exploit transactions. This discovery underscores Trace2Inv’s potential in unveiling new exploits.
@@ 7. Discussion
In this section, we discuss some key takeaways of this work and their implications in the context of invariant-based security measures for decentralized finance (DeFi) protocols.
@@
Choosing Complementary Invariants. Our findings underscore the importance of selecting a diverse set of invariants to safeguard smart contracts. Each type of invariant serves as a unique line of defense against abnormal transactions. For instance, invariants in time lock category act as temporal barriers, making it difficult for attackers to execute key functions like withdraw multiple times within one transaction. On the other hand, invariants in data and money flows categories limit the token amounts that can be withdrawn in one function call. By employing a combination of these invariants, developers force attackers to only withdraw a controlled amount of tokens per transaction, which may disrupt the mechanism the attack relies on, thereby blocking the attack.
@@
Dynamic Parameter Updates for Invariants. Another key insight is the need for dynamic parameter updates for certain types of invariants. For invariants that are tied to variables that change over time—like oracle prices or storage values—parameters can quickly become obsolete. If such an invariant is violated, it could lead to a cascade of failed transactions, causing a high FP rate. Thus, it is crucial for developers to monitor these variables and adjust the invariant parameters. Conversely, for invariants related to independent actions like token transfers, the parameters can remain relatively stable, as user behavior in these domains tends to be stable over time.
@@
Mitigating Flash Loan Attacks. Our benchmarks indicate a significant prevalence of flash loan-based hacks, with $17$ of $27$ examined hacks leveraging flash loan. Flash loans enable users to borrow large amounts of tokens for the duration of one transaction, providing attackers with substantial resources to execute complex hacks. Our approach can effectively mitigate the risk posed by flash loans. Invariants like EOA and OB block flash loan hacks by enforcing attackers to split their transaction into multiple ones. Without flash loan, the attacker would need to use their own assets to execute the hack with the uncertainty of other bots’ back-running between the hackers’ transactions. This raise not only the technical but also financial barriers to successful attacks.
@@
Impact of Invariants on Contract Composability. Incorporating invariant guards into smart contracts may limit their adaptability and integration with other DeFi protocols. However, our study in Section 4 reveals that many invariants stem from existing DeFi protocols requirements, underscoring the preference of developers for security benefits over flexibility. Moreover, our findings in RQ2, as discussed in Section 6, show that certain invariants, e.g., OB, have almost no FP, while others, e.g., TOU and DFU, can possibly block malicious transactions while allowing normal users to circumvent them. These insights imply that with careful selection and understanding of user behaviors, developers can devise invariant guards that minimally affect contract composability.
@@ 8. Related Work
Smart Contract Invariants. Prior works have studied various smart contract invariants. Zhou et al. (Zhou et al., 2020) introduces six invariants to defend against different hacks. Cider (Liu et al., 2022a) uses deep reinforcement learning to learn invariants that prevent arithmetic overflows, while SPCon (Liu et al., 2022b) recovers likely access control models from function callers in past transactions. Over (Deng et al., 2024) infers safety constraints on oracles using contract source code and oracle update history. In contrast, Trace2Inv explores a broader range of invariants and assesses their effectiveness and bypassability against real-world attacks.
@@
Smart Contract Security Analysis. There is a large body of works on the detection of security vulnerabilities in smart contracts. Oyente (Luu et al., 2016) is one of the first symbolic execution-based security tools to detect reentrancy, mishandled exception, transaction order dependency, and timestamp dependency. It has been extended to detect greedy, prodigal and suicidal contracts (Nikolić et al., 2018). Other well-known symbolic-execution tools include Manticore (man, 2019) and Mythril (myt, 2019) which are able to find other types of vulnerabilities, e.g., dangerous delegatecall, integer overflow, etc. Slither (sli, 2021) is another popular static security analysis tool for smart contracts. It performs data flow and control flow dependency analysis to support up to $87$ bug detectors. Other static analyzers include SmartCheck (Tikhomirov et al., 2018) mainly targeting bad coding practices, Securify (Sofware Reliability Lab, 2019) and Ethainter (Brent et al., 2020) for finding information-flow vulnerabilities. Moreover, dynamic analysis tools (Jiang et al., 2018; Atzei et al., 2017; Nguyen et al., 2020; Trail of Bits, 2019; Wüstholz and Christakis, 2020; Wang et al., 2020; Liu et al., 2020) were proposed to detect smart contract vulnerabilities through fuzzing and model-based testing.
@@
There is a substantial body of work focusing on the detection and exploitation of DeFi vulnerabilities. Qin et al. (Qin et al., 2021) and Zhou et al. (Zhou et al., 2021) manually formulated vulnerabilities such as oracle price manipulation and arbitrage into an optimization problem. FlashSyn (Chen et al., 2022) proposed a framework that automatically formulates flash loan attacks as a synthesis problem, by approximating the functions of DeFi protocols. Wu et al. (Wu et al., 2021) identified several oracle price manipulation patterns from on-chain transaction data to detect real-world attacks while Kong et al. (Kong et al., 2023) detects price manipulation vulnerabilities in DeFi applications through inter-contract taint analysis. Also, Gudgeon et al. (Gudgeon et al., 2020) showcased how to use flashloan to conduct governance attack. Baum et al. (Baum et al., 2022) surveyed the state-of-the-art mitigation techniques for front-running in DeFi. Interestingly, several works (Xue et al., 2022; Deng et al., 2023; Zhang et al., 2023) explored front-running as a defense mechanism against smart contract exploits. Different from the over-generalized security patterns used by the existing tools, our invariant guards capture the subtle semantic constraints of specific smart contracts.
@@
Runtime Verification and Validation. Runtime verification has been used for validation purpose where the runtime checks are on properties from users’ expectation rather than from formal program specifications (Magazzeni et al., 2017). Sereum (Rodler et al., 2018) is a general runtime validation framework to protect deployed contracts against reentrancy attacks. Solythesis (Li et al., 2020) provides a source-to-source compiler that compiles smart contracts with user-specified invariants to reject unexpected transactions (Li et al., 2020). In contrast, Trace2Inv is not limited to predefined attack types and has demonstrated effectiveness in mitigating a wide range of sophisticated attacks.